# Quantum Bootstrap Approach to a Non-Relativistic Potential for Quarkonium systems

### v1

@@ Abstract
I Introduction
The family of mesons known as quarkonium—bound states of a heavy quark and its own antiquark, such as charmonium ($c\bar{c}$) and bottomonium ($b\bar{b}$)—constitutes an exceptional laboratory for probing the theory of the strong interaction, Quantum Chromodynamics (QCD) [1, 2]. The large mass of the constituent quarks ensures that their motion within the bound state is largely non-relativistic, allowing their complex dynamics to be modeled by the non-relativistic Schrödinger equation with an effective potential [2]. This approach provides a crucial and tractable bridge between the full relativistic field theory of QCD and the rich, experimentally observed particle spectra.
@@ I Introduction
The application of this methodology to non-relativistic quantum mechanics is a more recent development[5, 7, 8, 9]. It relies on fundamental axiomatic principles—symmetry, consistency, and positivity—inherent to any valid quantum theory. These principles are translated into a set of algebraic recursion relations among expectation values of physical observables, combined with powerful positivity constraints on matrices of these moments [7, 8].This work applies the quantum bootstrap to the non-relativistic quarkonium potential, demonstrating its efficacy in a domain of central importance to particle physics.
The novelty of this approach lies in its non-perturbative robustness and computational efficiency. By leveraging fundamental consistency conditions, the bootstrap method provides a new pathway to determine quarkonium spectra with high precision. The results for the charmonium and bottomonium systems are presented, comparing them against experimental data from the Particle Data Group (PDG)[10].
@@
Furthermore, this analysis is extended to the frontier of the Standard Model: the toponium system. The top quark is unique; its extremely short lifetime ($\tau_{t}\approx 5\times 10^{-25}$ s) is an order of magnitude less than the hadronization timescale ($\tau_{\text{had}}\approx 3\times 10^{-24}$ s), precluding the formation of a stable bound state. However, recent experimental results from the CMS [11] and ATLAS [12] collaborations at the LHC have provided what could be evidence for a "quasi-bound" toponium state, manifesting as a cross-section enhancement near the $t\bar{t}$ production threshold. The theoretical bootstrap calculation for a hypothetical stable toponium provides a direct prediction for the mass of this observed resonance.
@@
This paper is organized as follows: Section II details the quantum bootstrap framework. Section III introduces the non-relativistic quarkonium potential and its physical basis in QCD. Section IV presents the numerical implementation and the results for charmonium and bottomonium spectra. Section V discusses the theoretical analysis of the hypothetical toponium system. Finally, Section VI offers some conclusions and perspectives for future research.
II The Quantum Bootstrap Framework
The quantum bootstrap method is constructed upon three fundamental principles: symmetry, consistency, and positivity.[5, 7, 8] It reframes the problem of finding energy eigenvalues from solving a differential equation to finding a consistent set of operator expectation values (moments) that satisfy a set of algebraic constraints.
Symmetry and Consistency: For any energy eigenstate $|\Psi\rangle$ with energy $E$ of a Hamiltonian $H$, the expectation values of operators must satisfy the identity $\langle\Psi|[H,\mathcal{O}]|\Psi\rangle=0$ for any operator $\mathcal{O}$ [4, 5]. This identity, a direct consequence of the time-independence of expectation values in an eigenstate, generates a network of linear recursion relations that connect moments of different orders. For a one-dimensional system with position operator $x$ and momentum $p$, choosing operators of the form $\mathcal{O}=x^{k}p^{m}$ allows one to systematically relate higher-order moments like $\langle x^{n}\rangle$ to lower-order ones. This algebraic structure reveals that the entire set of moments can be determined from a small set of fundamental data, which for the non-relativistic quarkonium potential includes the energy $E$ and a few low-order moments [5, 7].
Positivity: The cornerstone of the bootstrap method is the positivity of the norm in Hilbert space. This physical requirement translates into the powerful mathematical constraint that for any operator $\mathcal{O}$, the expectation value of its norm must be non-negative [4, 5]:
$\langle\Psi|\mathcal{O}^{\dagger}\mathcal{O}|\Psi\rangle\geq 0$ (1)
This constraint is made operational by constructing a "bootstrap matrix" of moments. For a one-dimensional radial problem, one can choose a basis of operators $\{1,r,r^{2},\dots,r^{K-1}\}$. The positivity condition implies that the $K\times K$ matrix $M$, known as a Hankel matrix [5], with elements
$M_{ij}=\langle r^{i+j}\rangle,\quad(i,j=0,\dots,K-1)$ (2)
must be positive semidefinite, i.e., all of its eigenvalues must be non-negative. The integer $K$ is referred to as the "depth" of the bootstrap, and increasing $K$ imposes more stringent constraints [5, 9].
The numerical algorithm thus involves scanning the space of the fundamental data (e.g., trial values for the energy $E$). For each point, the recursion relations are used to generate the required moments to populate the Hankel matrix. If the matrix is not positive semidefinite, that point is ruled out as physically inconsistent. The allowed solutions are confined to small "islands" in the parameter space, which converge rapidly to the true quantum spectrum as the matrix size $K$ is increased [5, 7, 8].
III Non-Relativistic Potential for Quarkonium Systems
A non-relativistic potential for quarkonium systems can be constructed to model the interaction between a heavy quark and its antiquark ($Q\bar{Q}$) [3]. Its functional form is a simple sum of a Coulomb-like term and a linear term, which captures the two essential features of the strong force as described by QCD :
$V(r)=-\frac{4}{3}\frac{\alpha_{s}}{r}+\kappa r$ (3)
Here, $r$ is the inter-quark separation, $\alpha_{s}$ is the strong coupling constant, and $\kappa$ is a parameter to be determined phenomenologically.
The first term, $\frac{4}{3}\frac{\alpha_{s}}{r}$, dominates at short distances ($r\to 0$) and represents the perturbative regime of QCD known as asymptotic freedom. It arises from the exchange of a single virtual gluon between the quark and antiquark, an interaction analogous to photon exchange in QED [2, 3]. The $1/r$ dependence is characteristic of a Coulomb-like force, and the prefactor of $4/3$ is the specific color factor for a $Q\bar{Q}$ pair in a color-singlet state.
The second term, $\kappa r$, dominates at large distances and models the non-perturbative phenomenon of color confinement [3]. Due to the self-interaction of gluons, the chromoelectric field lines between a distant quark and antiquark could be thought as collapsing into a narrow, one-dimensional flux tube [2, 3]. In this interpretation, the energy stored in this flux tube grows linearly with its length, giving rise to a potential that confines the quarks within hadrons, characterized by the parameter $\kappa$.
A key prediction of QCD in this interpretation is that this string tension should be flavor-independent, as it is a property of the vacuum itself, not the quarks. This has been confirmed by fits to both charmonium and bottomonium spectra, which yield consistent values for $\kappa$ [3]. The validity of this functional form for the potential is further substantiated by first-principles lattice QCD calculations of the static quark-antiquark potential, which reproduces the Coulomb-plus-linear shape [3].
IV Bootstrap Implementation and Results for Charmonium and Bottomonium
To apply the bootstrap method to quarkonium systems, the non-relativistic Schrödinger equation with the non-relativistic quarkonium potential is considered. For a state with orbital angular momentum $l$, the centrifugal barrier must be included, leading to the effective potential [2, 3]:
$V_{\text{eff}}(r)=\frac{\hbar^{2}l(l+1)}{2\mu r^{2}}-\frac{A}{r}+Br$ (4)
where $\mu$ is the reduced mass of the $Q\bar{Q}$ system, $A=\frac{4}{3}\alpha_{s}$, and $B=\kappa$. Applying the consistency condition $\langle[H,\mathcal{O}]\rangle=0$ with operators $\mathcal{O}$ built from powers of $r$ and $p$, a master recursion relation can be derived for the moments $\langle r^{m}\rangle$. For the non-relativistic quarkonium potential, this relation takes the form :
$\begin{split}(4m+2)B\langle r^{m}\rangle&=-4mE\langle r^{m-1}\rangle-(4m-2)A\langle r^{m-2}\rangle\\ &-\frac{\hbar^{2}}{2\mu}[m(m-1)(m-2)-4(m-1)l(l+1)]\langle r^{m-3}\rangle\end{split}$ (5)
The numerical solution begins by defining a two-dimensional search space spanned by the trial binding energy, $E$, and the first radial moment, $\textlangle r\textrangle$. While the normalization condition fixes the zeroth moment, $\textlangle r^{0}\textrangle=1$, the first moment, $\textlangle r\textrangle$, is not determined by the recursion relation for small values of m and thus serves as a necessary independent search parameter. For each point $(E,\textlangle r\textrangle)$ on the search grid, the entire tower of higher-order moments $\textlangle r^{m}\textrangle$} is uniquely determined by the iterative application of Equation (5). This equation is the computational core of this bootstrap algorithm. Given a trial energy $E$ and a few initial moments (which serve as the search parameters), this relation allows for the recursive generation of all higher-order moments needed to construct the Hankel matrix from Eq. 2. The algorithm proceeds with a direct grid search. For each node $(E,\textlangle r\textrangle)$ in the defined parameter space, the moments up to the order required to construct the $K\times K$ Hankel matrix are generated, where $K$ is the bootstrap depth. For this work, a depth of $K=5$ was chosen, requiring moments up to $\textlangle r^{8}\textrangle$. The positivity constraint is then tested by numerically computing the eigenvalues of this matrix. A point is deemed physically allowed only if all eigenvalues are non-negative, within a small numerical tolerance to account for floating-point inaccuracies. This process systematically maps out the regions of the parameter space consistent with the fundamental principles of quantum mechanics. The output of the grid search is a set of allowed points that form distinct, connected regions or ’islands’ in the $(E,\textlangle r\textrangle)$ plane. Each island corresponds to a specific physical bound state of the system. To determine the energy eigenvalue for a given state, the corresponding island is projected onto the energy axis. The resulting continuous band of allowed energies represents the solution at the given bootstrap depth. The final energy eigenvalue is estimated as the geometric center (midpoint) of this band. The width of the band provides a qualitative measure of the uncertainty, which systematically decreases as the bootstrap depth $K$ is increased.
This numerical procedure was implemented using reference parameters for the non-relativistic quarkonium potential and quark masses [3], which have been shown to sucessfully describe quarkonium spectroscopy. Specifically, $\alpha_{s}=0.2860$, $B=0.1306\text{ GeV}^{2}$, $m_{c}=1.488\text{ GeV}$ for charmoninum and $m_{b}=4.80303\text{ GeV}$, $B=0.18\text{ GeV}^{2}$ and $\alpha_{s}=0.39$ for bottomonium [3]. The masses for the ground S-wave ($l=0$) and P-wave ($l=1$) states of charmonium and bottomonium were calculated and compared with experimental values from the Particle Data Group (PDG) [10]. The results are summarized in Table 1.
@@ IV Bootstrap Implementation and Results for Charmonium and Bottomonium
The depth $K$ of the bootstrap controls the number of positivity constraints applied. While these results are presented for a fixed depth of $K=5$, a full analysis would involve studying the convergence of the predicted energy eigenvalues as $K$ is increased. As $K$ grows, the allowed ’islands’ shrink, and the energy predictions converge towards the exact values. This property, known as $K$-convergence, provides a powerful tool for quantifying and controlling the methodological uncertainty inherent in truncating the infinite set of positivity constraints. It is important to recognize that the truncation of the infinite set of positivity constraints at a finite $K$ introduces a methodological uncertainty. The bootstrap method has a property known as K-convergence, where the predicted energy eigenvalues stabilize as the matrix depth K is increased. A comprehensive uncertainty analysis would involve running the simulation for successively larger values of $K$ (e.g., $K=6,7$) to explicitly map this convergence and quantify the systematic error. While such a detailed scan is beyond the scope of this initial study, preliminary checks have been performed at $K=6$ which indicate that the $1S$ ground state masses for both charmonium and bottomonium shift by less than 1 MeV. This provides confidence that the results at $K=5$ are already close to the converged limit and that the methodological uncertainty from $K$-truncation is sub-dominant to the uncertainties inherent in the phenomenological potential model itself.

### v3

@@ Abstract
1 Introduction
The family of mesons known as quarkonium—bound states of a heavy quark and its own antiquark, such as charmonium ($c\bar{c}$) and bottomonium ($b\bar{b}$)—constitutes an exceptional laboratory for probing the theory of the strong interaction, Quantum Chromodynamics (QCD) [1, 2]. The large mass of the constituent quarks ensures that their motion within the bound state is largely non-relativistic, allowing their complex dynamics to be modeled by the non-relativistic Schrödinger equation with an effective potential [2]:
$V(r)=-\frac{4}{3}\frac{\alpha_{s}}{r}+\kappa r.$ (1)
This approach provides a crucial and tractable bridge between the full relativistic field theory of QCD and the rich, experimentally observed particle spectra.
@@ 1 Introduction
The application of this methodology to non-relativistic quantum mechanics is a more recent development[5, 7, 8, 9]. It relies on fundamental axiomatic principles—symmetry, consistency, and positivity—inherent to any valid quantum theory. These principles are translated into a set of algebraic recursion relations among expectation values of physical observables, combined with powerful positivity constraints on matrices of these moments [7, 8].This work applies the quantum bootstrap to the non-relativistic quarkonium potential, demonstrating its efficacy in a domain of central importance to particle physics. The novelty of this approach lies in its non-perturbative robustness and computational efficiency. By leveraging fundamental consistency conditions, the bootstrap method provides a new pathway to determine quarkonium spectra with high precision. The results for the charmonium and bottomonium systems are presented, comparing them against experimental data from the Particle Data Group (PDG)[10].
@@
Furthermore, this analysis is extended to the frontier of the Standard Model: the toponium system. The top quark is unique; its extremely short lifetime ($\tau_{t}\approx 5\times 10^{-25}$ s) is an order of magnitude less than the hadronization timescale ($\tau_{\text{had}}\approx 3\times 10^{-24}$ s), precluding the formation of a stable bound state. However, recent experimental results from the CMS [11] and ATLAS [12] collaborations at the LHC have provided what could be evidence for a ”quasi-bound” toponium state, manifesting as a cross-section enhancement near the $t\bar{t}$ production threshold. The theoretical bootstrap calculation for a hypothetical stable toponium provides a direct prediction for the mass of this observed resonance.
@@
This paper is organized as follows: Section 2 details the quantum bootstrap framework. Section 3 presents the details of the numerical implementation and section 4 presents the results for the proposed potential. Finally, Section 5 offers some conclusions and perspectives for future research.
2 Formalism and Moment Recursions
2.1 Quantum bootstrap framework
The quantum bootstrap reformulates the spectral problem of a Hamiltonian
$H=p^{2}+V(x)$ (2)
in terms of algebraic relations among operator moments $\mu_{n}=\langle x^{n}\rangle$ of an energy eigenstate ($H|{\psi}\rangle=E|{\psi}\rangle$). The expectation values of commutators and products of operators lead to recursion relations that depend only on the energy $E$ and the coefficients of the potential. It follows that:
$\langle[H,O]\rangle$ $=0,$ (3)
$\langle HO\rangle$ $=E\langle O\rangle.$ (4)
By choosing $O$ as powers of $x$ and $p$, one generates closed recursions among moments, without explicit reference to the wavefunction. In practice, one constructs the Hankel matrix
$M_{ij}=\langle x^{\,i+j}\rangle,\qquad i,j=0,1,\ldots,K-1,$ (5)
which must satisfy the positivity constraint $M\succeq 0$ because $\langle O^{\dagger}O\rangle=\mathbb{c}^{\dagger}M\mathbb{c}\geq 0$ for all polynomial operators $O=\sum_{i}c_{i}x^{i}$. Physical solutions correspond to energies $E$ and moment sets $\{\mu_{n}\}$ that simultaneously satisfy the recursion equations and all positivity constraints up to a chosen truncation depth $K$. As $K$ increases, the feasible intervals for $E$ contract exponentially around the exact eigenvalues.
2.2 Recursion relations for polynomial potentials
For a one-dimensional Hamiltonian $H=p^{2}+V(x)$ with $p=-i\partial_{x}$, the commutation relations $p^{2},x^{t}]=-2itp-(t(t-1))x^{t-2}$ yield the generic recursion
$4tE\,\langle x^{t-1}\rangle-4t\langle x^{t-1}V(x)\rangle+t(t-1)(t-2)\langle x^{t-3}\rangle-2\langle x^{t}V^{\prime}(x)\rangle=0.$ (6)
For polynomial potentials $V(x)=\sum_{k=0}^{d}a_{2k}x^{2k}$, Eq. (6) becomes a closed linear system connecting moments $\mu_{n}$ up to order n=t+2d1n=t+2d-1:
$\mu_{n+2d-1}=f_{n}(E,a_{2k})\,\mu_{n-2d+1}+\cdots,$ (7)
where the coefficients $f_{n}$ are polynomials in $E$ and the potential parameters. Solving these recursions allows one to express all higher moments in terms of a minimal dataset (e.g. $E$ and the lowest two even moments), which then become the optimization variables in the positivity constraints.
2.3 Radial formulation
For the study of quarkonium systems, we consider the radial Hamiltonian
$H_{r}=p_{r}^{2}+\frac{\ell(\ell+1)}{r^{2}}-\frac{A}{r}+Br,$ (8)
corresponding to the non-relativistic potential V(r)=A/r+BrV(r)=-A/r+Br, with $A=\frac{4}{3}\alpha_{s}$ and $B=\kappa$. In the bootstrap framework, expectation values are computed over radial wavefunctions $\psi(r)$ normalized on $r\in(0,\infty)$ with measure $r^{2}dr$. The appropriate operators are therefore constructed in terms of radial powers $r^{t}$ and the hermitian radial momentum operator
$p_{r}=-i\left(\frac{d}{dr}+\frac{1}{r}\right),$ (9)
which satisfies $p_{r},r]=-i$.
Applying the commutation identities (3)–(4) with $O=r^{t-1}$ and using
$p_{r}^{2},r^{t-1}]=-2i(t-1)p_{r}r^{t-2}-(t-1)(t-2)r^{t-3},$ (10)
Choosing $O=r^{m-1}$ and using the hermitian radial momentum $p_{r}=-i(\partial_{r}+1/r)$ with $p_{r},r]=-i$, one obtains after a straightforward but careful commutator algebra the closed recurrence for radial moments. In natural units ($\hbar=c=1$) it reads:
$(4t+2)B\,\langle r^{t}\rangle=-\,4t\,E\,\langle r^{t-1}\rangle\;-\,(4t-2)A\,\langle r^{t-2}\rangle\;-\;\frac{1}{2\mu}\Big[t(t-1)(t-2)-4(t-1)\ell(\ell+1)\Big]\langle r^{t-3}\rangle$ (11)
where $\mu$ is the reduced mass. Equation (11) expresses the moment $\langle r^{t}\rangle$ in terms of lower-order moments and the energy $E$. For a given orbital quantum number $\ell$, this recursion can be iterated up to the desired truncation order to construct the Hankel and auxiliary Stieltjes matrices that enforce positivity on the half-line $r\in(0,\infty)$.
2.4 Positivity on the half-line and auxiliary blocks
Unlike the full-line problem, where moments satisfy a single Hankel positivity condition, the bootstrap on $(0,\infty)$ requires two coupled positive semidefinite (PSD) constraints [9]:
$M^{(0)}_{ij}=\langle r^{i+j}\rangle\succeq 0,\qquad M^{(1)}_{ij}=\langle r^{i+j+1}\rangle\succeq 0.$ (12)
The first corresponds to the positivity of $\langle O^{\dagger}O\rangle$, while the second ensures that $\langle r\,O^{\dagger}O\rangle\geq 0$ for all polynomial operators $O(r)$. This pair of Hankel-type matrices defines the so-called Stieltjes moment problem, appropriate for radial domains. In numerical implementations, both PSD constraints are enforced simultaneously within a semidefinite programming (SDP) solver. Feasible pairs $(E,\{\mu_{n}\})$ satisfying Eqs. (11)–(18) define the allowed energy “islands” at truncation depth $K$.
2.5 Depth and convergence
The bootstrap depth $K$ controls the highest-order moments included in the positivity matrices. For polynomial potentials, the error in the extracted energy eigenvalues typically decays exponentially with $K$ [9, 8]. For the potential in equation 1, this convergence is accelerated by the smoothness of $V(r)$ and the absence of singularities for $\ell\geq 1$. As shown later in Sec. 4, accurate results for the lowest charmonium and bottomonium states are obtained already at moderate depths $K\lesssim 30$, corresponding to matrices of dimension $\sim 30\times 30$.
This formalism provides the theoretical foundation for the numerical bootstrap algorithm described in Sec. 3.
3 Numerical Algorithm and Implementation
The computational workflow proceeds in a hierarchy of constrained steps:
  1. 1.
    Initialization. Specify the potential $V(r)$, quantum number $\ell$, the reduced mass $\mu$, and fix a maximum bootstrap depth $K$.
  2. 2.
    Seed data. Choose a trial energy $E$ and a minimal set of low-order moments $\{\mu_{0},\mu_{1},\ldots,\mu_{m}\}$ compatible with normalization.
  3. 3.
    Moment generation. Use the recursion (11) to compute all higher-order $\mu_{n}$ up to order 2K22K-2.
  4. 4.
    Matrix assembly. Build the Hankel and auxiliary Stieltjes matrices $M^{(0)}$ and $M^{(1)}$ of dimensions $K\times K$ with entries $M^{(0)}_{ij}=\mu_{i+j}$ and $M^{(1)}_{ij}=\mu_{i+j+1}$.
  5. 5.
    Positivity test. Verify that both matrices are positive semidefinite. Energies satisfying $M^{(0)}\succeq 0$ and $M^{(1)}\succeq 0$ define the feasible set $\mathcal{S}_{K}$.
  6. 6.
    Refinement. Increase $K$ or apply semidefinite programming to narrow the energy interval $\mathcal{S}_{K+1}\subset\mathcal{S}_{K}$.
To accelerate convergence, the positivity test is recast as a convex optimization problem. For fixed $E$ the Hankel matrix is linear in the unknown low-order moments, $M^{(0)}(E,\mathbb{x})=\sum_{n=0}^{m}x_{n}F_{n}(E)$. Then the linear matrix inequality is solved:
$\max_{t,\mathbb{x}}\ t\quad\text{s.t.}\quad M^{(0)}(E,\mathbb{x})-tI\succeq 0,\qquad M^{(1)}(E,\mathbb{x})\succeq 0.$ (13)
The scalar $t_{\max}(E)$ measures the minimal eigenvalue margin. If $t_{\max}(E)>0$ the chosen $E$ belongs to the allowed set at depth $K$. Successive scans in $E$ and increasing $K$ yield exponentially tightening bounds on the true eigenvalues.
The algorithm was implemented in Python using the CVXPY interface to SDPA-GMP for arbitrary-precision SDP. All computations were performed with 50–100 bit precision, and matrices up to $K=30$ ($900\times 900$ elements) were feasible on a standard workstation. The total run time for a full energy scan of 200 points typically remains below one minute per $\ell$ sector. The algorithmic flow is shown in figure 1.
Figure 1: Algorithmic flow of the radial quantum bootstrap with SDP feasibility checks. The feasible energy set $\mathcal{S}_{K}$ contracts monotonically with depth $K$.
The convergence of the method is monitored by the shrinkage of $\mathcal{S}_{K}$ with $K$. For polynomial and smooth radial potentials, the midpoints of the intervals converge exponentially toward the true eigenvalues, $\delta E_{K}\propto e^{-\alpha K}$ with $\alpha\simeq 0.2$$0.3$. The precision limit is set by floating-point conditioning of the moment matrices; for large $K$, scaling the moments by $\langle r^{2}\rangle$ and enforcing normalization $\mu_{0}=1$ stabilizes the SDP.
This procedure yields energy estimates and moment sets that satisfy all algebraic and positivity constraints up to the chosen depth. These results form the basis for the physical analysis of quarkonium spectra presented in Sec. 4.
4 Results and Discussion
4.1 Bootstrap spectra for heavy quarkonia
We applied the bootstrap algorithm described in Sec. 3 to the potential in equation 1, using parameters from the literature ([13, 14, 15]), $\alpha_{s}=0.39$ and $\kappa=0.185~\text{GeV}^{2}$, together with reduced masses $\mu_{c}=1.48~\text{GeV}$, $\mu_{b}=4.72~\text{GeV}$, and $\mu_{t}=86.0~\text{GeV}$ for charmonium, bottomonium, and a hypothetical toponium, respectively. The bootstrap depth was typically $K=30$ for $\ell=0$ and $K=25$ for $\ell=1$ states, producing $30\times 30$ moment matrices and convergence to within $\sim 10^{-3}$ relative accuracy. The resulting spin-averaged centroid masses are summarized in Table 1.
@@ 4.1 Bootstrap spectra for heavy quarkonia
The depth $K$ of the bootstrap controls the number of positivity constraints applied. While these results are presented for a fixed depth of $K$, a full analysis would involve studying the convergence of the predicted energy eigenvalues as $K$ is increased. As $K$ grows, the allowed ’islands’ shrink, and the energy predictions converge towards the exact values. This property, known as $K$-convergence, provides a powerful tool for quantifying and controlling the methodological uncertainty inherent in truncating the infinite set of positivity constraints. It is important to recognize that the truncation of the infinite set of positivity constraints at a finite $K$ introduces a methodological uncertainty. The bootstrap method has a property known as $K$-convergence, where the predicted energy eigenvalues stabilize as the matrix depth $K$ is increased. A comprehensive uncertainty analysis would involve running the simulation for successively larger values of $K$ to explicitly map this convergence and quantify the systematic error. While such a detailed scan is beyond the scope of this initial study, preliminary checks indicate that the $1S$ ground state masses for both charmonium and bottomonium shift by a few MeV. This provides confidence that the results are already close to the converged limit and that the methodological uncertainty from $K$-truncation is sub-dominant to the uncertainties inherent in the phenomenological potential model itself.